4. Buffer capacity refers to the amount of acid or base a buffer can absorb without a significant change. It is governed by the concentrations of the acid and base components o conjugate buffer a d or base as a 0.1 M H the A can absorb times as much ac and concentration and given pH In this problem you begin with a bufer o known pH change. calculate the new pH after a particular quantity of acid or base is added. starting composition You are given 60 mL ofo.50 M phosphate buffer, 6.83, to test. The two major of the buffer, both in t of the concentration and the molar quantity of the phosphate species is: Molar quantity of HPO 8,2 mmol Concentration of HPO 0.304 M Concentration of HaPO. :0.196 M Molar quantity of HPO. :11.8 mmol a. You add 1.7 mL of 1.00 M HCI to the buffer. Calculate the molar quan ty of H,0 added as HCI, and the final molar quantity of HPO, and H,PO, at equilibrium? b. What is the new HPO 2-/H,PO ratio, and the new pH of the solution? The pKa of HPO. is 6.64. Use the Henderson-Hasselbach equation to calculate the new pH. (Note: You can use the molar ratio rather than the concentration ratio because both species are in the same volume.) c. Now take a fresh 60 mL of the 0.50 M pH 6.83 buffer and add 37 ml. of 1.00 M NaoH. Using steps similar to those above, calculate the new pH of the solution.
Answer #1

4. Buffer H2PO4-/HPO4^2-

a. molar quantity of H3O+ added = 1.0 M x 1.7 ml = 1.7 mmol

new molar quantity of HPO4^2- = 18.2 - 1.7 = 16.5 mmol

new molar quantity of H2PO4^- = 11.8 + 1.7 =13.5 mmol

b. New pH

using Hendersen-Hasselbalck equation,

pH = pKa + log(HPO4^2-/H2PO4-)

      = 6.64 + log(16.5/13.5)

      = 6.73

c. When,

3.7 ml of 1.0 M NaOH is added to buffer

molar quantity of OH- added = 1.0 M x 3.7 ml = 3.7 mmol

new molar quantity of HPO4^2- = 18.2 + 3.7 = 21.9 mmol

new molar quantity of H2PO4^- = 11.8 - 3.7 = 8.1 mmol

New pH

using Hendersen-Hasselbalck equation,

pH = pKa + log(HPO4^2-/H2PO4-)

      = 6.64 + log(21.9/8.1)

      = 7.072